# arXiv:2603.00415v1[eess.SY] 28 Feb 2026

### v1

@@ IV-B3 UPS-Integrated BESS
The UPS-integrated BESS, labeled as power smoothing units in Fig. I, extend this concept by supporting long-duration outage operation and enabling participation in demand response programs, while the UPS maintains continuity of critical loads when neither the grid nor the BESS can fully supply demand. Although this architecture increases installation cost, it can significantly improve overall system reliability [100].
@@ IV-C BESS Cost and Sizing Considerations
In [77], an evolutionary algorithm is used to generate a Pareto front across different battery capacities, enabling selection of an appropriate size based on the desired balance between operational benefits and battery degradation impacts. In [141], the BESS capacity is determined using a bi-level optimization framework in which the upper layer searches for the optimal storage size that minimizes the total system cost. For each candidate size, the lower layer performs optimal microgrid operation, including data center, and the BESS capacity that yields the lowest combined investment and operational cost is selected. On the other hand, [142] has investigated the BESS sizing for a data center based on reliability requirements rather than cost optimization. The required storage capacity is obtained by estimating the minimum battery backup time needed to maintain uninterrupted operation under stochastic grid outages and generator failures.
@@ VII Rack- and Server-Level ESSs roles in AI DCs
TABLE III: High-level comparison of rack-level BBUs and centralized UPS systems.
@@ VII-B1 BBU vs centralized UPS
BBUs and centralized UPS systems can both respond quickly in principle to supply the gap between grid power disturbance or outage and longer-term energy storage systems coming online (e.g., gensets / BESS), but they differ in where energy is buffered and what the buffer is optimized for. Centralized UPS concentrates energy and power conversion at room/facility scale, simplifying management and enabling grid-interactive use cases when permitted [35, 110]. Rack BBUs push storage behind fewer conversion stages in DC racks and reduce some double-conversion penalties compared to traditional centralized UPS architectures [115]. Practically, BBUs are best viewed as a distributed energy availability layer for the rack (seconds–minutes), not as the primary solution for sub-second smoothing which is handled more effectively by server/GPU capacitive buffers. Table III summarizes the high level comparison of rack-level BBUs and centralized UPS systems.
@@ VIII-B Hybrid GiUPS configuration for AI DCs using SLBESS
Figure 7: FFR by utilizing new BESS and SLBESS in GiUPS architecture
@@ IX Challenges and Future Directions
IX-A Challenges for Integrating ESSs in AI DCs
In the following section the main challenges and gaps are discussed considering multi-level ESSs in AI DCs. As discussed in previous sections, the ESSs not only can support the reliable operation of AI DCs but also can smooth the AI DC grid integration. In addition, it showed that integrating ESSs in different levels in AI DC power infrastructure leads to a grid interactive AI DC that can support the grid in emergency cases like frequency regulation, reactive power compensation or peak shaving scenarios. However, there are still challenges for implementing hierarchical ESSs in AI DCs, which will be discussed in the following sections.
IX-A1 Simulation and Experimental Validation
@@
To validate the optimal operation of the energy management system (EMS) of an AI DC, it is necessary to model its different components, including the ESSs. To our best knowledge, there is no software specifically developed for simulating EMSs of AI DCs. While MATLAB/Simulink or GT-SUITE software can be used to integrate mechanical and electrical parts in an AI DC, the integration of multiple solvers, electrical units, and mechanical units is not properly supported. Specifically, the multiple power conversion stages inside an AI DC require transient analysis, whereas the thermal units do not have the same requirement [13, 68]. Thus, accurate analysis of ESSs in AI DCs requires a software package that can simultaneously integrate transient and steady-state analyses [22]. On the other hand, the available real-time simulator devices are also not suitable for real-time experimental analysis of AI DCs, since they are primarily developed for analyzing electrical circuits rather than combined systems.
IX-A2 GPU Scheduling
@@
Current GPU-based AI workloads are predominantly executed at the highest supported core frequency, prioritizing peak performance at the expense of energy efficiency. Unlike CPUs, where dynamic voltage and frequency scaling (DVFS) is a mature and widely adopted mechanism, energy-aware frequency regulation for GPUs remains at an early stage of development [21, 138]. Moreover, existing GPU scheduling policies largely optimize for throughput and latency, without explicitly considering the energy consumption characteristics of heterogeneous deep learning tasks [19, 20]. This limitation is further exacerbated by the fact that many deep learning developers possess limited insight into the underlying hardware behavior of AI workloads. This can accelerate the degradation of ESSs and reveals the lack of control over GPU power profiles, which is essential for power smoothing scenarios.
IX-A3 Distribution Transformers and AI DCs
@@
Traditional distribution transformers are primarily designed for unidirectional power flow, supplying electricity from the medium-voltage grid to downstream loads, and their deployment in grids with grid-interactive AI DCs introduces new operational challenges. AI DCs equipped with large-scale ESSs and advanced power electronics can dynamically shift between high consumption and active power injection modes for grid services. Under such bidirectional power flow, conventional transformers may experience accelerated thermal aging due to loading patterns and reverse power conditions that deviate from standard design assumptions [75]. Moreover, voltage regulation mechanisms and tap settings optimized for passive load behavior may become ineffective or unstable when facing rapid transitions between load and injection states [43]. These limitations reveal that conventional distribution transformers are not inherently suited for highly dynamic, bidirectional operation driven by GiUPS or BESSs in AI DCs, underscoring the need for enhanced transformer models, real-time monitoring, and power-electronic-aware grid interfaces.
IX-A4 Load Forecasting in AI DCs
As discussed earlier, the AI DC power profile is hard for forecast. The traditional loads in power grids have predictable trajectories that can be predicted considering the proper historical data. However, this is not applicable for AI DC load profile. As a result, it can directly affected the sizing and ESS operation modes in AI DCs. For instance, the rack-level BESS or chip-level batteries which are designed for power smoothing applications need the accurate power consumption to operate effectively. In addition, the grid-scale BESSs are also need the accurate load profile of the DC to smooth the load and provide effective grid support.
IX-A5 Advanced ESSs Degradation Modeling and Life Time Prediction
@@
Due to high power variability in AI DCs, more frequent charging and discharging scenarios for ESSs are inevitable. This require a precise degradation analysis to maximize the life time and optimize the ESS operation. As discussed in section VII, the ESS at chip or rack level experiences a very high frequency in different operating conditions to smooth the power profile. In this regard, the online monitoring of the ESS could help for better life time prediction and proactive maintenance operations. As Lithium-ion batteries are more desirable for future AI DCs, the study of cyclic aging is important in both small scale ESSs or large scale BESSs. In addition, the installed SLBESSs in AI DCs are more vulnerable in case of aging and they need more precise monitorization especially for larger scale storage units.
IX-A6 Hierarchical ESSs Coordination
@@
Multi-level ESS implementation in AI DCs requires a centralized EMS to monitor the operating modes of different ESSs and ensure coordination among them [67]. The on-site grid-scale BESS should operate properly with the GiUPS to enable accurate power sharing, either to efficiently utilize RESs or to support the grid in emergency cases. Moreover, the GiUPS battery size and its switching time are related to the size of the on-site grid-scale BESS. Reliability analysis and cost-comparison analysis are required to determine the optimal combination of both for supplying the AI DC and supporting the power grid. The work [63] demonstrates a novel control framework for utilizing supercapacitors and grid-scale BESS for load smoothing, which shows the potential of hierarchical coordination among ESS units.
TES is also required to operate coherently with the on-site BESS or FC. The BESS absorbs excess power during off-peak periods, and this energy can be utilized to support the TES, as stated in section VI. However, coordinating these systems according to the BESS capacity and its key role in power smoothing or grid support is challenging. In addition, the BBUs and chip-level ESSs should operate together for power smoothing through multi-layer coordination. Thus, ESSs have a direct impact on each other, or at least on their neighboring layers, in a multi-layer structure, which necessitates a centralized control and management system with appropriate data acquisition units.
IX-A7 Long Duration ESSs (LDESSs)
On-site large scale BESSs or GiUPS systems require LDESSs to compensate the intermittency of RESs or providing long lasting AI DC support in case of the utility power failure [41]. The technology for LDESSs is still not mature and further investigation are needed.
IX-A8 Optimal Sizing and Cost Considerations
@@
There is little public research on the optimal sizing and cost analysis of ESSs inside DCs. In AI DCs, due to the presence of multi-layer ESSs, sizing and cost analysis are even more critical. For instance, is a grid-scale BESS more cost-effective, or a GiUPS system with high power-density batteries? Alternatively, is it more cost-effective to consider large-scale supercapacitors at the rack level, or to use a combination of BBUs and supercapacitors in load-smoothing scenarios? To answer these questions, multiple analyses should be considered to enable precise comparisons.
IX-B Possible Future Directions
IX-B1 Power-Aware GPU Scheduling
@@
During the training of deep learning models for predicting responses under different configuration settings, it is essential to jointly learn and optimize both throughput and energy consumption. This requires adaptive strategies for local and global batch size scheduling to balance convergence speed with computational efficiency [48]. Some solutions are discussed in section VII. In addition, GPU resources should be dynamically scheduled across training tasks by accounting for time-varying electricity prices and the availability of renewable energy sources, enabling energy-aware model training [16, 134, 62, 26]. To support such flexibility, elastic resource allocation is a key requirement, allowing the number of GPUs assigned to a training job to be adjusted dynamically during execution without disrupting the learning process. This elasticity enables efficient utilization of heterogeneous compute resources and provides the controllability of GPU power consumption which is necessary for chip-level ESS or BBU scheduling. Beyond this, dynamic batch size selection can be implemented to regulate the utility grid voltage profile [71]. The work [71] demonstrates the potential of chip-level utilization to enhance system-level objectives, in which ESSs implementation can also be included in the optimization problem.
IX-B2 AI DC Load Observability
@@
High-resolution monitoring of DC load dynamics can be achieved by employing waveform measurement units capable of capturing load profiles at very short time intervals (below 10 ms) [15]. Such fine-grained measurements enable the observation of fast transient behaviors and rapid load fluctuations that are otherwise invisible to conventional metering infrastructure. Building on this high-fidelity data, machine learning techniques can be leveraged to accurately predict short-term and long-term load patterns, supporting proactive control and optimization strategies. Besides, advanced load modeling studies are further increase the load observability of AI DCs not only for short term studies but also for long term planning and operation tasks [72, 85, 56, 83]. Furthermore, close collaboration between utilities and DC operators is essential to enable secure access to real-time load profiles, ensuring data confidentiality while facilitating coordinated grid–DC operation and enhanced system reliability.
IX-B3 Advanced Degradation, State Estimation, and RUL Prediction during operation
@@
Lithium-ion batteries undergo both calendar degradation and cyclic aging [65]. Calendar degradation is time-dependent and occurs regardless of battery usage, while cyclic degradation results from charge-discharge cycles. Degradation pathways, including solid electrolyte interphase (SEI) growth, lithium plating, and particle fracture [50] lead to capacity fade and increased internal resistance. To avoid the SLBs from entering the non-linear region of the degradation curve, a degradation-aware dispatch strategy needs to be devised. Estimating the RUL of SLBs is essential for safe deployment. Approaches range from data-driven techniques [28], neural network-based models, to entropy-based algorithms [122]. Uniquely, aging curves of SLBs are shaped by first-life usage patterns, making every battery distinct. European Union led Initiative Battery Passport [39] tracks lifecycle data, empowering refurbishers to select candidates for reuse.
IX-B4 FTM grid-scale BESS as A Reliable Reserve
@@
By monitoring and communicating the grid-scale BESS status, located in AI DC sites, with the grid operator, more optimized reserve planning can be achieved [140, 102]. In this case, the total amount of unused energy of grid-scale BESS can be considered a potential reserve that can be injected into the utility grid during contingency events such as the tripping of synchronous generators. The FTM grid-scale BESS can also play the role of an active reserve source in emergency cases by locally feeding the AI DC load. In this case, the grid operator disconnects the AI DC from the grid and allows the AI DC to supply its load using the available BESS co-generation resources. However, this requires precise coordination and control strategies between both the utility and AI DC owners.
IX-B5 Revisiting SSTs for AI DCs

### v2

@@ IV-B3 UPS-Integrated BESS
The UPS-integrated BESS, labeled as power smoothing units in Fig. 1, extend this concept by supporting long-duration outage operation and enabling participation in demand response programs, while the UPS maintains continuity of critical loads when neither the grid nor the BESS can fully supply demand. Although this architecture increases installation cost, it can significantly improve overall system reliability [98].
@@ VII Rack- and Server-Level ESSs roles in AI DCs
TABLE IV: High-level comparison of rack-level BBUs and centralized UPS systems.
@@ VII-B1 BBU vs centralized UPS
BBUs and centralized UPS systems can both respond quickly in principle to supply the gap between grid power disturbance or outage and longer-term energy storage systems coming online (e.g., gensets / BESS), but they differ in where energy is buffered and what the buffer is optimized for. Centralized UPS concentrates energy and power conversion at room/facility scale, simplifying management and enabling grid-interactive use cases when permitted [34, 108]. Rack BBUs push storage behind fewer conversion stages in DC racks and reduce some double-conversion penalties compared to traditional centralized UPS architectures [113]. Practically, BBUs are best viewed as a distributed energy availability layer for the rack (seconds–minutes), not as the primary solution for sub-second smoothing which is handled more effectively by server/GPU capacitive buffers. Table IV summarizes the high level comparison of rack-level BBUs and centralized UPS systems.
@@ VIII-B Hybrid GiUPS configuration for AI DCs using SLBESS
Figure 8: FFR by utilizing new BESS and SLBESS in GiUPS architecture
@@ IX Challenges and Future Directions
This section synthesizes the challenges and future research directions arising from the multi-layer ESS integration reviewed in this paper. We first address current challenges for both AI DC operation and ESS integration, then discuss future technical directions. Table V provides a consolidated cross-technology comparison.
TABLE V: Comparative Summary of Energy Storage Solutions for AI DCs
Technology Response Time Application Level Key Advantages Key Challenges Relative Cost AI DC Role Grid Interaction
Chip-level capacitors / HBM buffers Sub-ms to ms Chip Eliminates GPU power spikes; protects sensitive hardware Limited energy capacity; thermal management Low per-unit; high aggregate GPU power smoothing None (local)
Rack BBU / supercapacitors ms to seconds Rack/Server Fast transient response; modular deployment Limited energy; degradation under high-cycle loads Moderate Server backup; rack-level buffering None (local)
TUPS <20<20 ms Facility Proven reliability; load isolation Passive operation; idle batteries; no grid services Moderate–High Load ride-through Minimal
GiUPS 0.5–10 s (FFR) Facility Bidirectional operation; active grid support; FFR/VRT capability Complex control; regulatory approval; battery stress High FFR, VRT, peak shaving, frequency regulation Active (FFR, VRT, peak shaving)
BTM BESS Minutes–hours DC-scale Demand charge reduction; RES integration; backup support High CapEx; large footprint; sizing complexity High Load smoothing; backup; RES support Moderate (via PCC)
FTM BESS Minutes–hours Grid-scale Large capacity; ancillary services; reserve support Permitting; grid interconnection; high CapEx Very High Grid reserve; emergency supply Strong
Fuel Cell (FC) Seconds–minutes Facility Low-carbon operation (H2); high energy density; diesel replacement Fuel supply chain; cost; H2 infrastructure High Clean backup generation; co-generation Moderate
TES Minutes–hours Facility Cooling load shifting; RES utilization; no electrochemical degradation Integration complexity; auxiliary power; space requirement Moderate Cooling efficiency; peak load reduction Indirect
SLBESS Seconds–hours DC-scale / facility Lower CapEx; circular economy; scalable deployment SoH uncertainty; cell heterogeneity; safety standards Low–Moderate Cost-effective BTM/FTM backup Moderate
IX-A Current Challenges for AI Data Centers
IX-A1 AI DC Load Characteristics and Grid Impact
As discussed in the Introduction and Section II, AI DC load profiles differ fundamentally from traditional loads. The sub-second power variability of training workloads causes harmonic distortion, voltage flicker, and sub-frequency mechanical oscillations in synchronous generators [90, 99]. These impacts are amplified when multiple GPU clusters train synchronously. At the transmission level, the sudden switch of large AI DCs to backup power, as demonstrated by the 1,500 MW event in Virginia [79], reveals that current grid planning frameworks are insufficient for accommodating these highly responsive large loads. The primary challenge is the combination of high power magnitude, rapid and poorly predictable load transitions, and the absence of coordinated grid-response protocols between DC operators and utility grid operators.
IX-A2 Simulation and Experimental Validation
@@
To validate the optimal operation of the energy management system (EMS) of an AI DC, it is necessary to model its different components, including the ESSs. To the best of the authors’ knowledge, there is no software specifically developed for simulating EMSs of AI DCs. While MATLAB/Simulink or GT-SUITE can integrate mechanical and electrical components, the multi-physics nature of AI DCs, requiring simultaneous transient analysis for power electronics and thermal units alongside steady-state analysis for grid-scale systems, is not well supported by existing tools [12, 66, 21]. Available real-time simulator devices are similarly not suitable for real-time experimental analysis of AI DCs, since they are primarily developed for electrical circuits rather than combined electro-thermal-mechanical systems. This gap limits the ability to validate hierarchical ESS coordination strategies before physical deployment.
IX-A3 GPU Scheduling and Energy-Aware Compute
@@
Current GPU-based AI workloads are predominantly executed at the highest supported core frequency, prioritizing peak performance at the expense of energy efficiency. Unlike CPUs, where DVFS is a mature and widely adopted mechanism, energy-aware frequency regulation for GPUs remains at an early stage [20, 137]. Existing GPU scheduling policies largely optimize for throughput and latency without explicitly accounting for the energy consumption characteristics of heterogeneous deep learning tasks [18, 19]. The resulting lack of controllability over GPU power profiles directly constrains the effectiveness of chip-level and rack-level ESSs in power smoothing applications, since these systems must cope with power profiles that are neither observable nor controllable from the ESS perspective.
IX-A4 Distribution Transformers and Bidirectional Power Flow
@@
Traditional distribution transformers are designed for unidirectional power flow. AI DCs equipped with large-scale ESSs and advanced power electronics can dynamically shift between high consumption and active power injection for grid services. Under bidirectional power flow, conventional transformers experience accelerated thermal aging due to loading patterns that deviate from standard design assumptions [74]. Voltage regulation mechanisms and tap settings optimized for passive load behavior may become ineffective or unstable during rapid transitions between load and injection states [42]. These limitations underscore the need for enhanced transformer models, real-time monitoring, and power-electronic-aware grid interfaces at AI DC interconnection points.
IX-A5 Load Forecasting for ESS Sizing and Dispatch
Traditional grid loads have predictable trajectories that can be forecast using historical data. AI DC load profiles are fundamentally different: training job submissions, checkpoint saves, and idle periods create highly stochastic power demand that cannot be reliably modeled with conventional forecasting techniques. As a result, ESS sizing methodologies based on historical load patterns may under- or over-estimate required capacity. Rack-level BESSs and chip-level batteries designed for power smoothing require accurate sub-second power consumption data, while grid-scale BESSs require accurate load profiles at the facility level for effective grid support. Advanced machine learning-based load forecasting methods [71, 83, 55, 81] and high-resolution measurement infrastructure [14] are essential prerequisites for reliable ESS design and operation in AI DCs.
IX-A6 Advanced ESS Degradation Modeling and Lifetime Prediction
@@
Due to high power variability in AI DCs, more frequent and irregular charging and discharging cycles for ESSs are inevitable. This requires precise degradation analysis to maximize lifetime and optimize ESS operation. As discussed in Section VII, ESSs at the chip or rack level experience high-frequency cycling under varying conditions to smooth the power profile. Online monitoring can support better lifetime prediction and proactive maintenance. For lithium-ion batteries, cyclic aging is a key concern at both the small scale (BBU, chip-level) and large scale (BESS). SLBESSs are particularly vulnerable to aging effects and require more rigorous monitoring, especially for larger units where cell-level heterogeneity amplifies uncertainty in state-of-health estimation [64, 49].
IX-A7 Hierarchical ESS Coordination
@@
Multi-layer ESS implementation in AI DCs requires a centralized EMS to monitor and coordinate operating modes across all ESS layers [65]. The on-site grid-scale BESS must operate coherently with the GiUPS to enable accurate power sharing, RES utilization, and emergency grid support. The GiUPS battery size and switching time are fundamentally coupled to the size of the on-site BESS. This relationship requires combined reliability and cost analyses. The work [62] demonstrates a novel control framework utilizing supercapacitors and grid-scale BESS for load smoothing, illustrating the potential of hierarchical coordination. TES systems must also operate coherently with the BESS and FC. BBUs and chip-level ESSs must coordinate for multi-layer power smoothing. These interdependencies necessitate a centralized control and management system with appropriate data acquisition infrastructure spanning all layers.
IX-A8 Long-Duration Energy Storage (LDES)
On-site large-scale BESSs and GiUPS systems require long-duration energy storage (LDES) capability to compensate renewable intermittency or to sustain AI DC operation during extended utility outages [40]. Current LDES technologies are not yet mature for cost-effective deployment at the scale required by hyperscale AI DCs, and further investigation is needed into chemistries, system integration, and control strategies suitable for these applications.
IX-A9 Optimal Sizing and Cost Considerations
@@
There is limited public research on the optimal sizing and cost analysis of multi-layer ESSs in AI DCs. For instance, it is unclear whether a grid-scale BESS is more cost-effective than a GiUPS system with high power-density batteries, or whether large-scale supercapacitors at the rack level are preferable to a combination of BBUs and supercapacitors in load-smoothing scenarios. Answering these questions requires multi-objective techno-economic frameworks that integrate backup energy constraints, committed grid-service capacity, degradation-aware lifecycle cost modeling, and uncertainty in AI DC load profiles. Such frameworks are currently absent from the literature.
IX-B Future Directions for AI Data Centers and Energy Storage
IX-B1 Power-Aware GPU Scheduling and Chip-Level ESS Co-Design
@@
A fundamental open problem is the co-design of GPU execution strategies with chip-level and rack-level ESSs. During training, it is essential to jointly optimize throughput and energy consumption through adaptive strategies for local and global batch size scheduling [47]. GPU resources should be dynamically scheduled across training tasks by accounting for time-varying electricity prices and renewable energy availability, enabling energy-aware model training [15, 133, 61, 25]. Elastic resource allocation—dynamically adjusting the number of GPUs assigned to a training job—provides the controllability of GPU power consumption that is necessary for chip-level ESS or BBU scheduling. Beyond this, dynamic batch size selection can be implemented to regulate utility grid voltage profiles [70], directly linking compute scheduling to grid-level objectives. The integration of such power-aware scheduling with ESS dispatch optimization is an unexplored but promising research direction.
IX-B2 AI DC Load Observability and Forecasting Infrastructure
@@
High-resolution monitoring of DC load dynamics requires waveform measurement units capable of capturing load profiles at very short time intervals (below 10 ms) [14]. Such fine-grained measurements enable observation of fast transient behaviors that are invisible to conventional metering infrastructure. Building on high-fidelity data, machine learning techniques can predict short-term and long-term load patterns to support proactive ESS control and optimization [71, 83, 55, 81]. Advanced load modeling studies further increase the load observability of AI DCs for both short-term operation and long-term planning. Close collaboration between utilities and DC operators is essential to enable secure, real-time access to load profiles, supporting coordinated grid–DC operation.
IX-B3 Advanced Degradation Modeling, State Estimation, and RUL Prediction
@@
Lithium-ion batteries undergo both calendar and cyclic degradation [64]. Degradation pathways including SEI growth, lithium plating, and particle fracture [49] lead to capacity fade and increased internal resistance. For SLBs in particular, aging curves are shaped by first-life usage patterns, making each battery unique. Estimating the remaining useful life (RUL) is essential for safe deployment, with approaches ranging from data-driven techniques [27] and neural network-based models to entropy-based algorithms [120]. A degradation-aware dispatch strategy is needed to avoid pushing cells into the non-linear knee-point region. The EU Battery Passport initiative [38] provides a lifecycle data framework that can inform second-life deployment decisions. Extending such frameworks to AI DC operating conditions, with their high cycling frequency and variable load profiles, is an important open research direction.
IX-B4 FTM Grid-Scale BESS as a Reliable Reserve Asset
@@
By monitoring and communicating the grid-scale BESS status at AI DC sites with grid operators, more optimized reserve planning can be achieved [139, 100]. The total unused energy of the grid-scale BESS can be considered a potential reserve that can be injected into the grid during contingency events such as the tripping of synchronous generators. The FTM grid-scale BESS can also serve as an active reserve source in emergency cases by locally supplying the AI DC load in an islanded mode. This requires precise coordination and control strategies between utility and AI DC operators, as well as appropriate communication and dispatch protocols. Research on market mechanisms and regulatory frameworks that allow AI DCs to participate in reserve markets remains limited.
IX-B5 Revisiting Solid-State Transformers (SSTs) for AI DCs